System of Linear Equations and Problem Solving:
Can someone show me how to set with this word problem please?
An airplane traveled 645 mph with the wind and 575 mph against the wind. Find the speed of the plane in still air and the speed of the wind.

Answer by richwmiller(17219)


p+w=645
p-w=575
2p=1220
p=610 mph in still air
645-610=35 wind